What is the sequence in -9,101,-999,10001,-99999?
Answer by Edwin McCravy(20056)

-9,101,-999,10001,-99999?

    -9 = -10+1
   101 =  100+1
  -999 = -1000+1
 10001 =  10000+1
-99999 = -100000+1

So it's the powers of 10 with alternating signs, plus 1.

(-1)n is the factor that causes the signs to 
alternate:

So the nth term of the sequence is

an = (-1)n10n + 1

or

an = (-10)n + 1

Edwin